Is the ratio of the percentage by mole of elements in a compound equal to the ratio of no of atoms of the corresponding elements ?

Grade:12th Pass

3 Answers

Sunil Kumar FP
10 years ago
yes both are equal because there is a constant calles avagadro no whoch gets cancelled out.
let me explain it to u clearly
suppose the compound is MxNy
1mole of tye compound contains x mole of M and y mole of N
%composition=(x/x+y)×100 =M N=(y/x+y)×100=N

ratio=x/y which is equal ro their composition ratio.

thanks and regards
sunil kr
askIItian faculty
SANIKA
10 years ago
Will it be valid only for composition by mole or by mass also?
Sunil Kumar FP
10 years ago
No it will not be so. consider the example of Al2O3.let us take 102 gm of AL203 It implies 1 mole.therefore 1 mole of al203 will contain 2 mole of aluminiunm and 3 mope of oxygen.in mass % al=54/102×100 and mass percentage of oxygen=48/102×100. now ratio of their mass percentage =54/48 =9/8 which is not the ratio of 2:3. Thanks and regards sunil kumar askIItian faculty